OpenStudy (anonymous):

A line has a rise of 6 and a slope of 1/20. What is the run? 30 90 120

OpenStudy (freckles):

\[\text{ slope } =\frac{ \text{ rise } }{ \text{ run }}\]

OpenStudy (♪chibiterasu):

The rise is the numerator so we have to convert 1/20 so that the numerator matches.

OpenStudy (♪chibiterasu):

We do that by multiplying each part of the fraction by 6, since 1 * 6 = 6 and that would be the numerator.

OpenStudy (lynfran):

slope=rise/run 1/20=6/x x/20=6 x=6*20

OpenStudy (♪chibiterasu):

^

OpenStudy (anonymous):

thank you both!!